A total of $4000 was invested, part of it at 8% interest and the remainder at 9%. If the total yearly interest amounted to $350, how much was invested at each rate?
Answer by checkley77(12844)

.09x+.08(4,000-x)=350
.09x+320-.08x=350
.01x=350-320
.01x=30
x=30/.01
x=$3000 amount invested @ 9%
4,000-3000=1,000 amount invested @8%
Proof:
.09*3000+.08*1,000=350
270+80=350
350=350